heat of formation Fuel cells can be powered by the reaction between formic acid and oxygen as following HCOOH + 12 02 → CO2 + H2O As formic aci
Answer #1

The reaction occuring is

HCOOH(L) + \frac{1}{2}O_{2}(g)\rightarrow CO_{2}(g) + H_{2}O(g)

Formic acid is fed at 25°C and 1 atm

At T = 25°C formic acid is liquid

V = 100 L/min = 0.1 m3/min

T = 25°C = 298 K

Density of formic acid at T = 25°C = 1220 Kg/m3

Formic acid at inlet = (0.1) (1220) = 122 Kg/min

M. W of formic acid = 46 Kg/kmol

Moles of formic acid = 122/46 =

2.65217 kmol/min

formic acid is in outlet so all formic acid reacts

Air supllied is 100% excess

Oxygen required according to stiochiometry = (2.65217) (0.5) = 1.326085 moles

But o2 supplied = 1.326085(2) = 2.65217 mol/min

Air supplied = 2.65217/(0.21) =

12.6293 kgmol/min

The products leave at 70°C

Cp of products is taken at average temperature of (25+70) /2 = 47.5°C = 320.65 K

Reference enthaply is at T= 25°C

So ∆H(reactants) = 0

Product analysis

Component kmol/min Cp(KJ/mol°C) H(KJ/min)
CO2 2.65217 38.070 4543.565
H2O 2.65217 33.693 4021.18
O2 1.326085 29.536 762.52
N2 12.6293(0.79) = 9.9772 29.146 13085.796
Total 25.6029 22413.061

H(products) = 22413.061 KJ/min

From handbook at T = 25°C

∆Hf(CO2) (g) = -393. 5 KJ/mol

∆Hf(H2O) (g) = -241. 83 KJ/mol

∆Hf(HCOOH) (L) = -425. 5 KJ/mol

∆Hr = ∆Hproducts - ∆Hreactants

∆Hr = (-241.83) +(-393.5) -(-425.5) =

-209. 83 KJ/mol

Moles of formic acid reacted =2.65217 kmol/min

∆Hr = -209. 83(1000) (2.65217) =

-556504. 8311 KJ/min

Total energy

Q = ∆Hr + H(products)

Q = -556504. 8311 + 22413.061

Q = -534091. 7701 KJ/min = -8901. 529 KW

Fuel cell transfers to the surroundings 1.50(1000) KW of heat = 1500 KW

Doing energy balance electrical energy produced = 8901.529 - 1500 = 7401.529 KW

Electrical power = 7401.529 KW

Please upvote if helpful